Grade 11General Physics

The primary coil of an ideal step-up transformer has 200 turns and the transformation ratio is also 200. The input voltage the power are 220 V and 1100 W respectively. Calculate i) voltage across the secondary (ii) the current in the secondary ii) power in the secondary

tanisha singh
12 Years agoGrade 11

1 Answer

Saurabh Koranglekar
6 Years ago

Given data:
Primary coil turns, Np = 200
Transformation ratio, n = 200
Input voltage, Vp = 220 V
Input power, Pp = 1100 W

Step 1: Find Voltage Across the Secondary Coil
The transformation ratio is given by:
n = Ns / Np = Vs / Vp

Since n = 200, we substitute values:
200 = Vs / 220

Solving for Vs:
Vs = 200 × 220
Vs = 44000 V

Step 2: Find Current in the Secondary Coil
For an ideal transformer, power remains the same in both primary and secondary coils. Therefore,
Pp = Ps

Ps = Vs × Is
1100 = 44000 × Is

Solving for Is:
Is = 1100 / 44000
Is = 0.025 A

Step 3: Find Power in the Secondary Coil
Since the transformer is ideal, there is no power loss, so:
Ps = Pp = 1100 W

Final Answers:
(i) Voltage across the secondary coil, Vs = 44000 V
(ii) Current in the secondary coil, Is = 0.025 A
(iii) Power in the secondary coil, Ps = 1100 W